OpenStudy (anonymous):

Show that R={( a 0 / b c)} | a,b, c are elements of Z} i a ring with identity. a 0 / b c is the matrix with a 0 the first row and b c the second row.

OpenStudy (anonymous):

also show that the map f:R -> Z given by f(a 0 / b c) = a is a surjective homomorphism. What is the kernal of f?

terenzreignz (terenzreignz):

What does a 0 / b c mean?

OpenStudy (anonymous):

i didnt know how to make the matrix \[\left[\begin{matrix}a & 0 \\ b& c\end{matrix}\right]\]

terenzreignz (terenzreignz):

apparently, now you do :D

OpenStudy (anonymous):

well on the initial input math they dont give you the equation tools

terenzreignz (terenzreignz):

You have to memorise a few LaTeX tricks... but let's focus on your question, shall we? :)

OpenStudy (anonymous):

yes please

terenzreignz (terenzreignz):

You know how to multiply matrices?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

Well, then, for your first question, your only charge is pretty much to show closure... and show that matrix multiplication distributes over matrix addition... have you done this already?

OpenStudy (anonymous):

I can show closure but you'll have to elaborate on distributes over addition

terenzreignz (terenzreignz):

I mean, show that for 2x2 matrices A,B, and C A(B + C) = AB + AC (B + C)A = BA + BC

OpenStudy (anonymous):

oh show distribution okay

OpenStudy (anonymous):

that proves identity though?

terenzreignz (terenzreignz):

That's one of the ring-axioms, right? :D

OpenStudy (anonymous):

okay okay we are going to prove its a ring and then prove it has identity. I was going too far

terenzreignz (terenzreignz):

So you have it from here?

OpenStudy (anonymous):

except also show that the map f:R -> Z given by f(a 0 / b c) = a is a surjective homomorphism. What is the kernal of f?

terenzreignz (terenzreignz):

Is it a homomorphism, though? Show that first...

terenzreignz (terenzreignz):

show that for any two matrices A and B in R f(A + B) = f(A) + f(B) and f(AB) = f(A)f(B) That should be enough to show homomorphism-ness :D

OpenStudy (anonymous):

how do i determine the kernel?

terenzreignz (terenzreignz):

Kernel is simply the subset of R which is mapped to 0... namely, those 2x2 matrices which have 0 for their upper-left entry :)

OpenStudy (anonymous):

oh yeah thanks! I should be able to do it from here. I think I was just trying to over complicate it.

terenzreignz (terenzreignz):

No problem :)

OpenStudy (anonymous):

wait! how to I know its surjective? I mean to me it looks obvious that it is an onto function but is there a formal way to show this?